# Effective Learning with Node Perturbation in Multi-Layer Neural Networks

### v4

@@ Appendix D Algorithm pseudocode
        for layer $l$ from $1$ to $L$ do {Regular forward pass}
@@
        for layer $l$ from $1$ to $L$ do {Noisy forward pass}
@@
           $\mathbf{\tilde{a}}_{l}=\mathbf{W}_{l}\mathbf{\tilde{x}}^{\star}_{l-1}+\bm{\epsilon}_{l}$
@@
        $\delta\mathcal{L}=(\mathbf{t}-\mathbf{\tilde{x}}_{L})^{2}-(\mathbf{t}-\mathbf{x}_{L})^{2}$Compute loss difference}
@@
           $\mathbf{W}_{l}\leftarrow\mathbf{W}_{l}-\eta N\,\delta\mathcal{L}\,\frac{\mathbf{\tilde{a}}_{l}-\mathbf{a}_{l}}{||\delta\mathbf{{a}}||^{2}}\left({\mathbf{x}^{\star}_{l-1}}\right)^{\top}$Update weight matrix}
           $\mathbf{R}_{l}\leftarrow\mathbf{R}_{l}-\alpha\left(\mathbf{x}^{\star}_{l}\left(\mathbf{x}^{\star}_{l}\right)^{\top}-\textrm{diag}\left(\left(\mathbf{x}^{\star}_{l}\right)^{2}\right)\right)\mathbf{R}_{l}$Update decorrelation matrix}
@@
Appendix E Illustration of learning rules
Figure 5 illustrates the difference between different learning rules.
@@ Appendix F Experimental details
Experiments were performed using fully-connected or convolutional neural networks with leaky ReLU activation functions and a mean squared error (MSE) loss on the one-hot encoding of class membership. All experiments were run on an HP OMEN GT13-0695nd desktop computer with an NVIDIA RTX 3090 GPU. The experiments required about a week of total computation time and several more weeks of computation time during exploratory experiments.
@@
Table 1: The neural network architectures consist of fully connected (FC) and convolutional (Conv) layers. All layers except the output layer are followed by a leaky ReLU transformation.
@@
FC 100
The learning rates were determined for each algorithm separately by a grid search, in which the learning rate started at $10^{-6}$ and was doubled until performance stopped increasing. Table 2 shows the learning rates $\eta$ used in each experiment. For each experiment, the highest stable learning rate was selected. For the decorrelation learning rate a fixed value of $\alpha=10^{-3}$ was chosen based on a prior manual exploration. Note that the minibatch size was fixed for every method at 1000. When using NP methods, activity perturbations were drawn from a univariate Gaussian with variance $\sigma^{2}=10^{-6}$.
@@
Method 1 layer 3 hidden layers 6 hidden layers 9 hidden layers ConvNet
NP $8.2\times 10^{0}$ $2.6\times 10^{-1}$
DNP $1.0\times 10^{3}$ $1.3\times 10^{2}$
ANP $2.6\times 10^{-1}$ $1.6\times 10^{1}$
DANP $1.3\times 10^{2}$ $6.6\times 10^{1}$ $6.6\times 10^{1}$ $5.2\times 10^{2}$
INP $1.0\times 10^{-3}$ $6.4\times 10^{-2}$
DINP $2.6\times 10^{-1}$ $2.6\times 10^{-1}$ $2.6\times 10^{-1}$ $2.0\times 10^{0}$
BP $8.0\times 10^{-6}$ $6.4\times 10^{-5}$ $5.1\times 10^{-4}$
DBP $1.0\times 10^{-3}$ $4.1\times 10^{-3}$ $1.0\times 10^{-3}$
Appendix G Impact of perturbation scale on performance
@@
Appendix H Peak accuracies
Table 3: Peak percentage accuracies for different learning algorithms and architectures. Fully connected models consisting of 1, 3, 6 and 9 hidden layers were trained on CIFAR-10. ConvNets were trained on CIFAR-100. Best performance across all methods is shown in boldface.
@@ Appendix H Peak accuracies
NP 41.6 38.4 39.2 39.2 5.5 5.8
DNP 48.2 37.5 99.0 46.2 38.6 18.6
ANP 39.6 39.7 14.2 11.8
DANP 94.6 45.7 95.5 41.6 92.1 40.6 77.1 22.2
INP 44.4 43.2 21.1 16.6
DINP 99.9 47.8 100.0 47.3 100.0 46.5 84.6 24.7
BP 42.0 38.0 96.7 55.7 36.1 26.1
DBP 49.0 39.0 100.0 54.1 99.2 $\bm{29.7}$
Appendix I SARCOS experiment
Figure 7 shows train and test loss for (D)NP, (D)ANP and (D)BP for the SARCOS dataset [Vijayakumar and Schaal, 2000]. The task is to solve inverse dynamics problem for a seven degrees-of-freedom SARCOS anthropomorphic robot arm by mapping from a 21-dimensional input space (7 joint positions, 7 joint velocities, 7 joint accelerations) to the corresponding 7 joint torques. The network is a fully connected architecture with two 32-unit hidden layers and was trained with MSE loss and the leaky ReLU activation function. The network was trained for 500 epochs and results were averaged over three random seeds. A learning rate search for each algorithm was started at $10^{-7}$, doubling the learning rate until the algorithm became unstable. A decorrelation learning rate of $10^{-4}$ was used for DNP and DANP and $10^{-3}$ for DBP.

### v5

@@ Appendix E Algorithm pseudocode
         for layer $l$ from $1$ to $L$ do $\triangleright$ Regular forward pass
@@
         for layer $l$ from $1$ to $L$ do $\triangleright$ Noisy forward pass
@@
             $\mathbf{\tilde{a}}_{l}=\mathbf{W}_{l}\mathbf{\tilde{x}}^{\star}_{l-1}+{\bf\it\epsilon}_{l}$
@@
         $\delta\mathcal{L}=(\mathbf{t}-\mathbf{\tilde{x}}_{L})^{2}-(\mathbf{t}-\mathbf{x}_{L})^{2}$triangleright Compute loss difference
@@
             $\mathbf{W}_{l}\leftarrow\mathbf{W}_{l}-\eta N\,\delta\mathcal{L}\,\frac{\mathbf{\tilde{a}}_{l}-\mathbf{a}_{l}}{||\delta\mathbf{{a}}||^{2}}\left({\mathbf{x}^{\star}_{l-1}}\right)^{\top}$triangleright Update weight matrix
             $\mathbf{R}_{l}\leftarrow\mathbf{R}_{l}-\alpha\left(\mathbf{x}^{\star}_{l}\left(\mathbf{x}^{\star}_{l}\right)^{\top}-\textrm{diag}\left(\left(\mathbf{x}^{\star}_{l}\right)^{2}\right)\right)\mathbf{R}_{l}$triangleright Update decorrelation matrix
@@ Appendix F Experimental details
Experiments were performed using fully-connected or convolutional neural networks with leaky ReLU activation functions and a categorical cross-entropy (CCE) loss on the one-hot encoding of class membership. The Adam optimizer with default parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-7$, was used for optimization. All experiments were run on an HP OMEN GT13-0695nd desktop computer with an NVIDIA RTX 3090 GPU. The experiments required about a week of total computation time and several more weeks of computation time during exploratory experiments.
@@
Table 1: The neural network architectures consist of fully connected (FC) and convolutional (Conv) layers. All layers except the output layer are followed by a leaky ReLU transformation. Convolutional layer size is given as height×$\times$×width×$\times$×output channels, stride.
@@
FC 10
ConvNet Tiny ImageNet Conv 5$\times$5$\times$32, 2
Conv 3$\times$3$\times$64, 2
Conv 5$\times$5$\times$128, 1
FC 512
FC 512
FC 200
Table 2 shows the learning rates $\eta$ used in each experiment. For each experiment, the highest stable learning rate was selected. For the decorrelation learning rate a fixed value of $\alpha=10^{-3}$ was chosen based on a prior manual exploration. Note that the minibatch size was fixed for every method at 1000. When using NP methods, activity perturbations were drawn from a univariate Gaussian with variance $\sigma^{2}=10^{-6}$.
@@
Method 1 layer 3 hidden layers 6 hidden layers 9 hidden layers ConvNet ConvNet Tiny ImageNet
NP $1.0\times 10^{-3}$ $1.0\times 10^{-4}$
DNP $1.0\times 10^{-3}$ $1.0\times 10^{-3}$ $5.0\times 10^{-3}$ $5.0\times 10^{-3}$ $1.0\times 10^{-3}$
ANP $1.0\times 10^{-4}$ $1.0\times 10^{-4}$
DANP $1.0\times 10^{-3}$ $5.0\times 10^{-3}$ $5.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-3}$
INP $1.0\times 10^{-4}$ $1.0\times 10^{-4}$
DINP $1.0\times 10^{-3}$ $5.0\times 10^{-3}$ $5.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-3}$
BP $1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-3}$ $1.0\times 10^{-4}$
DBP $1.0\times 10^{-3}$ $1.0\times 10^{-3}$ $1.0\times 10^{-3}$
Appendix G Alignment as a function of noisy forward passes
Figure 6 shows the same data as Figure 2, but the $x$-axis displays the number of noisy forward passes performed, instead of the number of noise iterations. This figure essentially compares NP, ANP and INP given the same amount of computation.
Figure 6: Angles between BP’s weight update and an update calculated by NP, ANP and INP as a function of the number of noisy forward passes.
Appendix H Impact of perturbation scale on performance
@@
Appendix I Peak accuracies
Table LABEL:tab:peakacc shows peak accuracies for all algorithms used in the CIFAR-10 experiments and Table LABEL:tab:peakacc_tin shows peak accuracies for the Tiny ImageNet experiment.
Table 3: Peak percentage accuracies for different learning algorithms and architectures on CIFAR-10. Best performance across all methods is shown in boldface.
@@ Appendix I Peak accuracies
BP 41.3 39.3 59.6 51.5 100.0 62.1
DBP 44.9 40.9 99.6 54.3 100.0 60.5
NP 38.5 38.2 28.1 28.1 45.2 44.5
DNP 42.8 40.2 44.9 42.1 46.7 43.1 40.5 38.6 50.2 47.9
ANP 30.4 30.6 52.7 49.9
DANP 48.5 44.4 47.3 44.2 44.8 42.7 58.2 52.6
INP 32.0 32.0 52.3 49.9
DINP 52.6 46.3 53.0 46.6 52.0 45.3 57.9 52.2
Table 4: Peak percentage accuracies for different learning algorithms on Tiny ImageNet. Best performance across all methods is shown in boldface.
Tiny ImageNet
Method Train Test
BP 55.5 14.5
DANP 1 iter 2.2 2.1
DANP 10 iters 13.5 9.0
DANP 30 iters 32.8 11.3
DANP 100 iters 35.4 12.0
DINP 1 iter 2.0 2.0
DINP 10 iters 11.4 8.1
DINP 30 iters 26.7 11.8
DINP 100 iters 60.8 13.4
DNP 100 iters 9.7 7.5
Appendix J SARCOS experiment
Figure 8 shows train and test loss for (D)NP, (D)ANP and (D)BP for the SARCOS dataset (Vijayakumar & Schaal, 2000). The task is to solve inverse dynamics problem for a seven degrees-of-freedom SARCOS anthropomorphic robot arm by mapping from a 21-dimensional input space (7 joint positions, 7 joint velocities, 7 joint accelerations) to the corresponding 7 joint torques. The network is a fully connected architecture with two 32-unit hidden layers and was trained with MSE loss and the leaky ReLU activation function. The network was trained for 500 epochs and results were averaged over three random seeds. A learning rate search for each algorithm was started at $10^{-7}$, doubling the learning rate until the algorithm became unstable. A decorrelation learning rate of $10^{-4}$ was used for DNP and DANP and $10^{-3}$ for DBP.
@@
Appendix K Deeper CIFAR architectures
This section describes the scaling of DNP, DANP and DINP to deeper fully connected networks as well as convolutional neural networks.
Figure 9 (Top panel) compares the performance of DNP, DANP and DINP for 3, 6 and 9 hidden layers. Performance for DINP and DANP scales better with network depth than performance for DNP, again highlighting the improvement these novel formulations provide.
Figure 9 (Bottom panel) shows that, when training a convolutional neural network on CIFAR-10, DANP and DINP outperform NP, though relative to BP, a large gap in train accuracy emerges. Also, the NP-style algorithms lag behind BP more than they did in the fully-connected networks. All decorrelated NP variants massively outperform their non-decorrelated counterparts in both train and test accuracy. Interestingly, (D)ANP outperforms (D)NP in this experiment as well.
Figure 9: Performance comparisons. Curves report mean train and test accuracy. Shaded areas indicate the maximal and minimal accuracy obtained for three random seeds. Top: Performance of DNP, DANP and DINP for 3-, 6- and 9-hidden-layer networks. Bottom: Performance of different node perturbation and backpropagation variants when training a convolutional neural network on CIFAR-10.